OpenStudy (anonymous):

find x and y using the elimination method6x + 5y = 4 and 7x – 10y = -8

OpenStudy (anonymous):

7x-10y = -8 6x+ 5y = 4 Comparing the two equations this way, it's a little easier to see which variable we want to eliminate first. Note that in the first equation, we have -10y and in the second we have 5y and that 5*2=10. So let's multiply the bottom equation by 2: 7x -10y = -8 12x+10y = 8 Now, you can see that if we add the two equations, we can eliminate the +-10ys, giving us 19x = 0 x = 0 Now, just plug back in and solve for y!